# Methods to achieve near-millisecond energy relaxation and dephasing times for a superconducting transmon qubit

### v1

@@
Abstract
Superconducting qubits are one of the most promising physical systems for implementing a quantum computer. However, executing quantum algorithms of practical computational advantage requires further improvements in the fidelities of qubit operations, which are currently limited by the energy relaxation and dephasing times of the qubits. Here, we report our measurement results of a high-coherence transmon qubit with energy relaxation and echo dephasing times surpassing those in the existing literature. We measure a qubit frequency of 2.890 GHz, an energy relaxation time with a median of µs and a maximum of µs, and an echo dephasing time with a median of µs and a maximum of µs. We report details of our design, fabrication process, and measurement setup to facilitate the reproduction and wide adoption of high-coherence transmon qubits in the academia and industry.
NOTE ADDED TO arXiv VERSION: We appreciate any early feedback on the manuscript, which we will take into account before the journal submission planned for mid-August 2024.
@@ I Introduction
In the past two decades, the rapid development of superconducting qubits has rendered them one of the most promising candidates for realizing large-scale quantum computers [1, 2, 3, 4, 5]. The transmon qubit, proposed in 2007, has become the most widely used superconducting qubit due to its simplicity and performance [6, 7]. It consists of a Josephson junction and a shunt capacitor which reduces the sensitivity of the qubit to charge noise but keeps the qubit sufficiently anharmonic for realizing fast, high-fidelity gate operations.
Short coherence times have historically been a disadvantage of superconducting qubits, but the previous decade has shown continuous improvements on this challenge, leading to increased fidelities of qubit operations. The longest reported energy relaxation times of transmons have approached but not surpassed 500-µs median value as shown in Table 1 based on Refs. [8, 9, 10, 11, 12, 13, 14]. Recently, there have also been significant improvements in transmon echo dephasing times with the longest reported values extending to 300–400 µs [12, 9, 14]. With other types of superconducting qubits such as 0- qubits and fluxonium qubits, energy relaxation times above a millisecond have been observed, but these qubits have either had short dephasing times or low qubit frequencies [15, 16].
| Ref. | (GHz) | (µs) | (106) | (µs) |
| 8 | 3.73 | 303 | 7.1 | 201 |
| 9 | 3.92 | 401 | 9.9 | N/A |
| 10 | 3.86 | 393 | 9.5 | N/A |
| 11 | 4.24 | 292 | 7.7 | N/A |
| 12 | 3.02 | 270 | 5.1 | 307 |
| 13 | 3.95 | 323 | 8.0 | 286 |
| 14 | 4.79 | 210 | 6.5 | 290 |
| This work | 2.89 | 502 | 9.1 | 541 |
@@ II Sample and its fabrication
Figure 1 shows a microscope image of a sample which is identical to the one measured in this work. It contains four transmon qubits –, each of which couples to a coplanar waveguide resonator for readout. Qubits Q1 and Q3 are flux-tunable by the use of a superconducting quantum interference device (SQUID), whereas Q2 and Q4 are fixed-frequency qubits. The four qubit–resonator pairs are identical except for the size of the Josephson junction and the length of the resonator. The readout resonators couple to a shared Purcell filter, which reduces the energy relaxation of the qubits into the readout lines [17, 18]. The layout design file of the sample is available in Ref. [19].
This section provides a detailed description of our sample fabrication process, which is an adaptation of the recipe described in Ref. [14] to the chemicals and equipment available to us. Notable differences include the choice of CF4 as the processing gas for Nb etching, which improves the reproducibility of the etch, and the pre-dicing of the sample between the Al evaporation and lift-off, which minimizes the exposure of the sample to the ambient atmosphere. The pieces of equipment used in our process are listed in Table 2.
@@ II.0.1 Substrate and niobium patterning
The sample is fabricated using a 675-µm-thick 6-inch (100)-oriented high-resistivity (10 kcm) intrinsic-silicon wafer sourced from Siegert Wafer. The pre-cleaning of the wafer follows Ref. [20] and begins with an RCA solvent clean. The wafer then undergoes a dip in dilute hydrofluoric (HF) acid (1:50) for 1 min. After being rapidly transferred to the sputtering tool to minimize exposure to the ambient atmosphere, the wafer is baked at 300 ∘C under vacuum. The sputtering of a 200-nm Nb film is carried out near room temperature and at 2600-W power. The sputter target has a purity of 99.998%. After the sputtering, the wafer is coated with a protective layer of AZ 5214E photoresist and diced into 25 mm 30 mm rectangular coupons with a dicing saw.
@@ II.0.4 Dicing and liftoff
III Setup for qubit measurements
Our experimental setup used for qubit characterization and measurements is presented in Fig. 2, with detailed information on measurement equipment and components provided in Table 3. The sample is wire-bonded to a QCage.24 sample holder and placed inside QCage Magnetic Shielding, both of which are supplied by QDevil (under Quantum Machines). The sample is exposed to ambient atmosphere at room temperature for a total of 7 days. It is then cooled down to approximately 10 mK in a Bluefors dilution refrigerator equipped with a tin-plated copper shield at 10 mK and an Amumetal magnetic shield just inside the outer vacuum chamber (OVC). We directly generate the qubit control and readout signals without any analog mixers by using a Xilinx RFSoC evaluation board with QICK firmware [24, 25] locked to a Rb frequency standard. The signals then pass through rf attenuators, an Eccosorb infrared filter, and a low pass filter before reaching the sample. The readout signal coming out of the sample goes through a three-wave-mixing Josephson traveling-wave parametric amplifier (TWPA), a cryogenic high-electron-mobility-transistor (HEMT) amplifier, and a room-temperature HEMT amplifier before being digitized by RFSoC. During the first cooldown, the TWPA is not pumped and therefore does not provide any amplification.
@@ III Setup for qubit measurements
IV Measurement Results
| Qubit | (GHz) | (µs) | (µs) | (106) |
| (tunable) | 4.047 | 195 | 237 | 5.0 |
| (fixed) | 2.890 | 502 | 549 | 9.1 |
| (tunable) | 4.593 | 44 | 28 | 1.3 |
| (fixed) | 3.295 | 154 | 114 | 3.2 |
During the first cooldown, we measure that the fixed-frequency qubit exhibits especially long energy relaxation and echo dephasing times: medians of µs and of µs and maxima of µs and µs. The data for these results are shown in Fig. 3 in the form of distributions for repeated measurements of and and time traces for the maximum values. In each time trace, we repeat and average 500 measurements for each value of delay time between a -pulse and a readout pulse. The number of repetitions is for the measurement. The measurement results for all four qubits during the first cooldown is summarized in Table 4.
During the second cooldown, we carry out a more extensive measurement of the qubit . The qubit frequency is measured to have decreased by MHz compared to the first cooldown. The readout resonator of this qubit is measured to have a linewidth of 2.63 MHz and a dispersive shift of MHz. Then, we measure the energy relaxation and echo dephasing times with and without the pump signal applied to the TWPA. We fit an exponential function to the obtained data points, each of which is an average of 500 measurements, and compute the relative uncertainties of the fit parameters. We find that the TWPA does not significantly affect the coherence times but improves the certainties of the and obtained from the fit. These results are presented in Fig. 4. We exclude from further analysis the values with relative uncertainties exceeding and values with relative uncertainties exceeding . The resulting data are shown in Fig. 6. We obtain medians of µs and µs and maxima of µs and µs. Figure 5 shows the time traces for the maximum values.
V Discussion
In conclusion, this result represents a significant step in the development of high-coherence superconducting qubits by approaching the millisecond mark for the energy relaxation and dephasing times of a transmon qubit. It also demonstrates the robustness and reproducibility of the fabrication recipe described in Ref. 14 and the effectiveness of the equipment and components listed in Tables 2 and 3. Detailed reporting of a high-coherence qubit will benefit the research community and accelerate the global efforts on developing quantum sensors, quantum simulators, and quantum computers based on superconducting quantum technologies.
Acknowledgements.
We acknowledge the Research Council of Finland under its Centre of Excellence Quantum Technology Finland (Grant nos. 336810, 352925, 352934 and 352935), the European Research Council under the Advanced Grant ConceptQ (Grant no. 101053801), the Finnish Quantum Institute InstituteQ, and the provision of facilities and technical support by Aalto University at the OtaNano-Micronova Nanofabrication Centre.References
- Huang et al. [2020] H.-L. Huang, D. Wu, D. Fan, and X. Zhu, Superconducting Quantum Computing: A Review, Science China Information Sciences 63, 180501 (2020), arXiv:2006.10433 [quant-ph].
- Tuokkola et al. [2024] M. Tuokkola, Y. Sunada, H. Kivijärvi, L. Grönberg, J.-P. Kaikkonen, V. Vesterinen, J. Govenius, and M. Möttönen, Data and GDS file for ”Methods to achieve near-millisecond energy relaxation and dephasing times for a superconducting transmon qubit (2024).
Appendix A Extended Data Figures

### v2

@@
Methods to achieve near-millisecond energy relaxation and dephasing times for a superconducting transmon qubit
Tiivistelmä
Superconducting qubits are one of the most promising physical systems for implementing quantum computers. However, executing quantum algorithms of practical computational advantage requires further improvements in the fidelities of qubit operations, which are currently limited by the energy relaxation and dephasing times of the qubits. Here, we report our measurement results of a high-coherence transmon qubit with energy relaxation and echo dephasing times surpassing those in the existing literature. We measure a qubit frequency of 2.890 GHz, an energy relaxation time with a median of µs and a maximum of µs, and an echo dephasing time with a median of µs and a maximum of µs. We report in detail our design, fabrication process, and measurement setup to facilitate the reproduction and wide adoption of high-coherence transmon qubits in the academia and industry.
@@ I Introduction
In the past two decades, the rapid development of superconducting qubits has rendered them one of the most promising candidates for realizing large-scale quantum computers [1, 2, 3, 4, 5, 6]. The transmon qubit, proposed in 2007, has become the most widely used superconducting qubit due to its simplicity and performance [7, 8]. It consists of a Josephson junction and a shunt capacitor which reduces the sensitivity of the qubit to charge noise but keeps the qubit sufficiently anharmonic for realizing fast, high-fidelity gate operations.
Short coherence times have historically been a disadvantage of superconducting qubits, but the previous decade has shown continuous improvements on this challenge, leading to increased fidelities of qubit operations. The longest reported energy relaxation times of transmons have approached but not surpassed 500-µs median value as shown in Fig. 1 and Extended Data Table 1 based on Refs. [9, 10, 11, 12, 13, 14, 15, 16]. Recently, there have also been significant improvements in transmon echo dephasing times with the longest reported average extending to 307 µs [10, 13, 14, 16]. Transmon qubits with coherence times above 100 µs have also been reported, for example, in Refs. [17, 18, 19, 20, 21, 22]. With other types of superconducting qubits such as 0- qubits and fluxonium qubits, energy relaxation times above a millisecond have been observed, but these qubits have either had short dephasing times or low qubit frequencies [23, 24, 25].
@@ II Sample and its fabrication
Figure 2 shows a microscope image of a sample which is identical to the one measured in this work. It contains four transmon qubits –, each of which couples to a coplanar waveguide resonator for readout. Qubits Q1 and Q3 are flux-tunable by the use of a superconducting quantum interference device (SQUID), whereas Q2 and Q4 are fixed-frequency qubits. The four qubit–resonator pairs are identical except for the size of the Josephson junction and the length of the resonator. The readout resonators couple to a shared Purcell filter, which reduces the energy relaxation of the qubits into the readout lines [26, 27]. The layout design file of the sample is available in Ref. [28]. For more details of the sample, see Extended Data Figs. 1 and 2
This section provides a detailed description of our sample fabrication process, which is an adaptation of the recipe described in Ref. [16] to the chemicals and equipment available to us. Notable differences include the choice of CF4 as the processing gas for Nb etching, which improves the reproducibility of the etch, and the pre-dicing of the sample between the Al evaporation and lift-off, which minimizes the exposure of the sample to the ambient atmosphere. The pieces of equipment used in our process are listed in Extended Data Table 2.
@@ II.0.1 Substrate and niobium patterning
The sample is fabricated using a 675-µm-thick 6-inch (100)-oriented high-resistivity (10 kcm) intrinsic-silicon wafer sourced from Siegert Wafer. The pre-cleaning of the wafer follows Ref. [29] and begins with an RCA solvent clean. The wafer then undergoes a dip in dilute hydrofluoric (HF) acid (1:100) for 1 min. After being rapidly transferred to the sputtering tool to minimize exposure to the ambient atmosphere, the wafer is baked at 300 ∘C under vacuum. The sputtering of a 200-nm Nb film is carried out near room temperature and at 2600-W power. The sputter target has a purity of 99.998%. After the sputtering, the wafer is coated with a protective layer of AZ 5214E photoresist and diced into 25 mm 30 mm rectangular coupons with a dicing saw.
@@ II.0.4 Dicing and liftoff
III Setup for qubit measurements
Our experimental setup used for qubit characterization and measurements is presented in Fig. 3, with detailed information on measurement equipment and components provided in Extended Data Table 3. The sample is wire-bonded to a QCage.24 sample holder and placed inside QCage Magnetic Shielding, both of which are supplied by QDevil (under Quantum Machines). The sample is exposed to ambient atmosphere at room temperature for a total of 7 days. It is then cooled down to approximately 10 mK in a Bluefors dilution refrigerator equipped with a tin-plated copper shield at 10 mK and an Amumetal magnetic shield just inside the outer vacuum chamber (OVC). We directly generate the qubit control and readout signals without any analog mixers by using a Xilinx RFSoC evaluation board with QICK firmware [33, 34] locked to a Rb frequency standard. The signals then pass through rf attenuators, an Eccosorb infrared filter, and a low pass filter before reaching the sample. The readout signal coming out of the sample goes through a three-wave-mixing Josephson traveling-wave parametric amplifier (TWPA), a cryogenic high-electron-mobility-transistor (HEMT) amplifier, and a room-temperature HEMT amplifier before being digitized by RFSoC. During the first cooldown, the TWPA is not pumped and therefore does not provide any amplification.
| Qubit | (GHz) | (µs) | (µs) | (106) |
| (tunable) | 4.047 | 195 | 237 | 5.0 |
| (fixed) | 2.890 | 502 | 549 | 9.1 |
| (tunable) | 4.593 | 44 | 28 | 1.3 |
| (fixed) | 3.295 | 154 | 114 | 3.2 |
IV Measurement Results
During the first cooldown, we measure that the fixed-frequency qubit exhibits especially long energy relaxation and echo dephasing times: medians of µs and of µs and maxima of µs and µs. The data for these results are shown in Fig. 4(a, b, e, f) in the form of distributions for repeated measurements of and and time traces for the maximum values. In each time trace, we repeat and average 500 measurements for each value of delay time between a -pulse and a readout pulse. The number of repetitions is for the measurement. The measurement results for all four qubits during the first cooldown is summarized in Table 1.
During the second cooldown, we carry out a more extensive measurement of the qubit . The qubit frequency is measured to have decreased by MHz compared to the first cooldown. The readout resonator of this qubit is measured to have a linewidth of 2.63 MHz and a dispersive shift of MHz. Then, we measure the energy relaxation and echo dephasing times with and without the pump signal applied to the TWPA. We fit an exponential function to the obtained data points, each of which is an average of 500 measurements, and compute the relative uncertainties of the fit parameters. We find that the TWPA does not significantly affect the coherence times but improves the certainties of the and obtained from the fit. These results are presented in Extended Data Fig. 3. We exclude from further analysis the values with relative uncertainties exceeding and values with relative uncertainties exceeding . The resulting data are shown in Fig. 4(c, d, g, h) and Extended Data Fig. 4. We obtain medians of µs and µs and maxima of µs and µs. The time trace for the longest measured deviates from an exponential function, likely because of the qubit frequency shifting during the sweep owing to the charge noise.
V Discussion
In conclusion, this result represents a significant step in the development of high-coherence superconducting qubits by approaching the millisecond mark for the energy relaxation and dephasing times of a transmon qubit. It also demonstrates the robustness and reproducibility of the fabrication recipe described in Ref. 16 and the effectiveness of the equipment and components listed in Extended Data Tables 2 and 3. Detailed reporting of a high-coherence qubit will benefit the research community and accelerate the global efforts on developing quantum sensors, quantum simulators, and quantum computers based on superconducting quantum technologies.
Data availability
The data that support the findings of this study are available at https://doi.org/10.5281/zenodo.12819934.
Acknowledgements.
We acknowledge the Research Council of Finland under its Centre of Excellence Quantum Technology Finland (Grant nos. 336810, 352925, 352934 and 352935) and the Finnish Quantum Flagship, the European Research Council under the Advanced Grant ConceptQ (no. 101053801), the Finnish Quantum Institute InstituteQ, Business Finland through Quantum Technologies Industrial project (no. 41419/31/2020), Horizon Europe programme HORIZON-CL4-2022-QUANTUM-01-SGA through OpenSuperQPlus100 project (no. 101113946), Jane and Aatos Erkko Foundation, QDOC Doctoral Pilot Programme, and the provision of facilities and technical support by Aalto University at the OtaNano-Micronova Nanofabrication Centre.Competing interests
M.M. declares that he is a Co-Founder and Shareholder of the quantum-computer company IQM Finland and of the quantum-algorithm company QMill Oy. Other authors declare no competing interests.
Viitteet
- Huang et al. [2020] H.-L. Huang, D. Wu, D. Fan, and X. Zhu, Superconducting quantum computing: a review, Science China Information Sciences 63, 180501 (2020).
- Acharya et al. [2024] R. Acharya, L. Aghababaie-Beni, I. Aleiner, T. I. Andersen, M. Ansmann, F. Arute, K. Arya, A. Asfaw, N. Astrakhantsev, J. Atalaya, R. Babbush, D. Bacon, B. Ballard, J. C. Bardin, J. Bausch, A. Bengtsson, A. Bilmes, S. Blackwell, S. Boixo, G. Bortoli, A. Bourassa, J. Bovaird, L. Brill, M. Broughton, D. A. Browne, B. Buchea, B. B. Buckley, D. A. Buell, T. Burger, B. Burkett, N. Bushnell, A. Cabrera, J. Campero, H.-S. Chang, Y. Chen, Z. Chen, B. Chiaro, D. Chik, C. Chou, J. Claes, A. Y. Cleland, J. Cogan, R. Collins, P. Conner, W. Courtney, A. L. Crook, B. Curtin, S. Das, A. Davies, L. De Lorenzo, D. M. Debroy, S. Demura, M. Devoret, A. Di Paolo, P. Donohoe, I. Drozdov, A. Dunsworth, C. Earle, T. Edlich, A. Eickbusch, A. M. Elbag, M. Elzouka, C. Erickson, L. Faoro, E. Farhi, V. S. Ferreira, L. F. Burgos, E. Forati, A. G. Fowler, B. Foxen, S. Ganjam, G. Garcia, R. Gasca, E. Genois, W. Giang, C. Gidney, D. Gilboa, R. Gosula, A. G. Dau, D. Graumann, A. Greene, J. A. Gross, S. Habegger, J. Hall, M. C. Hamilton, M. Hansen, M. P. Harrigan, S. D. Harrington, F. J. H. Heras, S. Heslin, P. Heu, O. Higgott, G. Hill, J. Hilton, G. Holland, S. Hong, H.-Y. Huang, A. Huff, W. J. Huggins, L. B. Ioffe, S. V. Isakov, J. Iveland, E. Jeffrey, Z. Jiang, C. Jones, S. Jordan, C. Joshi, P. Juhas, D. Kafri, H. Kang, A. H. Karamlou, K. Kechedzhi, J. Kelly, T. Khaire, T. Khattar, M. Khezri, S. Kim, P. V. Klimov, A. R. Klots, B. Kobrin, P. Kohli, A. N. Korotkov, F. Kostritsa, R. Kothari, B. Kozlovskii, J. M. Kreikebaum, V. D. Kurilovich, N. Lacroix, D. Landhuis, T. Lange-Dei, B. W. Langley, P. Laptev, K.-M. Lau, L. L. Guevel, J. Ledford, K. Lee, Y. D. Lensky, S. Leon, B. J. Lester, W. Y. Li, Y. Li, A. T. Lill, W. Liu, W. P. Livingston, A. Locharla, E. Lucero, D. Lundahl, A. Lunt, S. Madhuk, F. D. Malone, A. Maloney, S. Mandrá, L. S. Martin, S. Martin, O. Martin, C. Maxfield, J. R. McClean, M. McEwen, S. Meeks, A. Megrant, X. Mi, K. C. Miao, A. Mieszala, R. Molavi, S. Molina, S. Montazeri, A. Morvan, R. Movassagh, W. Mruczkiewicz, O. Naaman, M. Neeley, C. Neill, A. Nersisyan, H. Neven, M. Newman, J. H. Ng, A. Nguyen, M. Nguyen, C.-H. Ni, T. E. O’Brien, W. D. Oliver, A. Opremcak, K. Ottosson, A. Petukhov, A. Pizzuto, J. Platt, R. Potter, O. Pritchard, L. P. Pryadko, C. Quintana, G. Ramachandran, M. J. Reagor, D. M. Rhodes, G. Roberts, E. Rosenberg, E. Rosenfeld, P. Roushan, N. C. Rubin, N. Saei, D. Sank, K. Sankaragomathi, K. J. Satzinger, H. F. Schurkus, C. Schuster, A. W. Senior, M. J. Shearn, A. Shorter, N. Shutty, V. Shvarts, S. Singh, V. Sivak, J. Skruzny, S. Small, V. Smelyanskiy, W. C. Smith, R. D. Somma, S. Springer, G. Sterling, D. Strain, J. Suchard, A. Szasz, A. Sztein, D. Thor, A. Torres, M. M. Torunbalci, A. Vaishnav, J. Vargas, S. Vdovichev, G. Vidal, B. Villalonga, C. V. Heidweiller, S. Waltman, S. X. Wang, B. Ware, K. Weber, T. White, K. Wong, B. W. K. Woo, C. Xing, Z. J. Yao, P. Yeh, B. Ying, J. Yoo, N. Yosri, G. Young, A. Zalcman, Y. Zhang, N. Zhu, and N. Zobrist, Quantum error correction below the surface code threshold (2024), arXiv:2408.13687 [quant-ph].
- Sivak et al. [2023] V. V. Sivak, A. Eickbusch, B. Royer, S. Singh, I. Tsioutsios, S. Ganjam, A. Miano, B. L. Brock, A. Z. Ding, L. Frunzio, S. M. Girvin, R. J. Schoelkopf, and M. H. Devoret, Real-time quantum error correction beyond break-even, Nature 616, 50 (2023).
- Dunsworth et al. [2017] A. Dunsworth, A. Megrant, C. Quintana, Z. Chen, R. Barends, B. Burkett, B. Foxen, Y. Chen, B. Chiaro, A. Fowler, R. Graff, E. Jeffrey, J. Kelly, E. Lucero, J. Y. Mutus, M. Neeley, C. Neill, P. Roushan, D. Sank, A. Vainsencher, J. Wenner, T. C. White, and J. M. Martinis, Characterization and reduction of capacitive loss induced by sub-micron Josephson junction fabrication in superconducting qubits, Applied Physics Letters 111, 022601 (2017).
- Wang et al. [2019] Z. Wang, S. Shankar, Z. Minev, P. Campagne-Ibarcq, A. Narla, and M. Devoret, Cavity Attenuators for Superconducting Qubits, Physical Review Applied 11, 014031 (2019).
- Nersisyan et al. [2019] A. Nersisyan, E. A. Sete, S. Stanwyck, A. Bestwick, M. Reagor, S. Poletto, N. Alidoust, R. Manenti, R. Renzas, C.-V. Bui, K. Vu, T. Whyland, and Y. Mohan, Manufacturing low dissipation superconducting quantum processors, in 2019 IEEE International Electron Devices Meeting (IEDM) (IEEE, San Francisco, CA, USA, 2019) pp. 31.1.1–31.1.4.
- Wei et al. [2020] K. X. Wei, I. Lauer, S. Srinivasan, N. Sundaresan, D. T. McClure, D. Toyli, D. C. McKay, J. M. Gambetta, and S. Sheldon, Verifying multipartite entangled Greenberger-Horne-Zeilinger states via multiple quantum coherences, Physical Review A 101, 032343 (2020).
- Moskalev et al. [2023] D. O. Moskalev, E. V. Zikiy, A. A. Pishchimova, D. A. Ezenkova, N. S. Smirnov, A. I. Ivanov, N. D. Korshakov, and I. A. Rodionov, Optimization of shadow evaporation and oxidation for reproducible quantum Josephson junction circuits, Scientific Reports 13, 4174 (2023).
- Pishchimova et al. [2023] A. A. Pishchimova, N. S. Smirnov, D. A. Ezenkova, E. A. Krivko, E. V. Zikiy, D. O. Moskalev, A. I. Ivanov, N. D. Korshakov, and I. A. Rodionov, Improving Josephson junction reproducibility for superconducting quantum circuits: junction area fluctuation, Scientific Reports 13, 6772 (2023).
- Pop et al. [2014] I. M. Pop, K. Geerlings, G. Catelani, R. J. Schoelkopf, L. I. Glazman, and M. H. Devoret, Coherent suppression of electromagnetic dissipation due to superconducting quasiparticles, Nature 508, 369 (2014).
- Tuokkola et al. [2024] M. Tuokkola, Y. Sunada, H. Kivijärvi, J. Albanese, L. Grönberg, J.-P. Kaikkonen, V. Vesterinen, J. Govenius, and M. Möttönen, Data and GDS file for ”Methods to achieve near-millisecond energy relaxation and dephasing times for a superconducting transmon qubit”missing (2024).
Liite A Extended Data Figures and Tables
| Ref. | (GHz) | (µs) | (106) | (µs) |
| 9 | 3.73 | 303 | 7.1 | 201 |
| 10 | 3.92 | 401 | 9.9 | N/A |
| 11 | 3.86 | 393 | 9.5 | N/A |
| 12 | 4.24 | 292 | 7.7 | N/A |
| 13 | 5.92 | 280 | 10.4 | 238 |
| 14 | 3.02 | 270 | 5.1 | 307 |
| 15 | 3.95 | 323 | 8.0 | 286 |
| 16 | 4.79 | 210 | 6.5 | 290 |
| This work | 2.89 | 502 | 9.1 | 541 |